OpenStudy (anonymous):

Prove that f(0) = 0.

OpenStudy (anonymous):

for what function ?

OpenStudy (anonymous):

f(x + y) = f(x) + f(y) from R to R

OpenStudy (anonymous):

any more info on this function. Maybe like it is eaven?

OpenStudy (anonymous):

f is defined from the reals to the reals for all x,y in the reals

OpenStudy (anonymous):

f(0+y)=f(0)+f(y) or f(y)=f(0)+f(y) so f(0) have to equal 0